OpenStudy (anonymous):

what is 15 and 7/10ths minus 12 and 2/5ths?

OpenStudy (anonymous):

2/5 = 4/10 by multiplying it by 2. so, 15 - 12 = 3, and 7/10 - 4/10 = 3/10. Answer: 3 and 3/10